OpenStudy (anonymous):

You have two exponential functions. One function has the formula g(x) = 3(2x). The other function has the formula h(x) = 2x+1. Which option below gives formula for k(x) = (g – h)(x)? k(x) = 2x k(x) = 5(2x) k(x) = 5(2x+1) k(x) = 2

OpenStudy (mathstudent55):

Are the functions \(g(x) = 3(2^x)\) and \(h(x) = 2^{x + 1} \) ? You need to write equations clearly, so we don't have to waste time trying to guess what they are. The second function can also be \(h(x) = 2^x + 1\) How do we know what you mean? Use the ^ (the caret symbol) to mean an exponent and use parentheses where needed.

OpenStudy (mathstudent55):

You can also use the equation editor, the draw tool, or post a picture.

OpenStudy (anonymous):

You have two exponential functions. One function has the formula g(x)=3(2^x). The other function has the formula h(x)=2^x+1. Which option below gives the formula for k(x)=(g-h)(x) a) k(x)=2^x b) k(x)=5(2^x) c) k(x)=5(2^x+1) d) K(x)=2

OpenStudy (jdoe0001):

hmmm should be a 2 there so hmm keep in mind that \((g-h)(x)\iff g(x)-h(x)\qquad \begin{cases} g(x)=3(2^x)\\ h(x)=2^x+1 \end{cases} \\ \quad \\ g(x)-h(x)\implies [3(2^x)]\quad -\quad [2^x+1]\)